Two objects undergo an elastic collision. Object 1
Answer #1

given kinetic energy of the object 1 is 51J then we know that kinetic energy is

a)

K.E=0.5mv^2

then velocity of the object 1 is

V1=sqrt((2*K.E)/m1)

=sqrt((2*51)/2.75)

=6.09m/s

b)

By the conservation of energy

K.E of initial system is equal to the final kinetic energy of the system

then kinetic energy of the cart system after the collision is equal to the 51J

c)

for this by the consertation of the momentum

momentum before the collision is equal to the momentum after the collision\

then initial momentum is

M1U1+M2U2

=2.75*6.09+2.6*0

=16.74kgm/s

then final momentum is equal to the 16.74kgm/s